OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (whovianchick):

−3|2x + 6| = −12 −3|2x + 6|/-3= −12/-3 |2x + 6| =4 2x+6=4 2x+6=-4 2x+6-6=4-6 2x+6-6=-4-6 2x=-2 2x=-10 x=-1 x=-5

OpenStudy (anonymous):

it b

OpenStudy (whovianchick):

Correct

OpenStudy (anonymous):

:P

OpenStudy (whovianchick):

I'm faster than this but the work takes forever to type out

OpenStudy (anonymous):

@whovianchick this is the only time i giv u a medal